# Edit Once, Update Everywhere: A Simple Framework for Cross-Lingual Knowledge Synchronization in LLMs

### v1

@@ Abstract
Knowledge editing allows for efficiently adapting large language models (LLMs) to new information or corrections without requiring full retraining. However, prior methods typically focus on single-language or basic multilingual editing, failing to achieve true cross-linguistic knowledge synchronization. To address this, we present a simple and practical state-of-the-art (SOTA) recipe Cross-Lingual Knowledge Democracy Edit (X-KDE), designed to propagate knowledge from a dominant language to other languages effectively. Our X-KDE comprises two stages: (i) Cross-lingual Edition Instruction Tuning (XE-IT), which fine-tunes the model on a curated parallel dataset to modify in-scope knowledge while preserving unrelated information, and (ii) Target-language Preference Optimization (TL-PO), which applies advanced optimization techniques to ensure consistency across languages, fostering the transfer of updates. Additionally, we contribute a high-quality, cross-lingual dataset, specifically designed to enhance knowledge transfer across languages. Extensive experiments on the Bi-ZsRE and MzsRE benchmarks show that X-KDE significantly enhances cross-lingual performance, achieving an average improvement of +8.19%, while maintaining high accuracy in monolingual settings.
@@
Yuchen Wu1, Liang Ding2thanks: Correspond to Liang Ding liangding.liam@gmail.com, Li Shen3, Dacheng Tao4 1Shanghai Jiao Tong University 2The University of Sydney 3Sun Yatsen University 4Nanyang Technological University
1 Introduction
Large Language Models (LLMs) Achiam et al. (2023); Dubey et al. (2024); Yang et al. (2024a); Guo et al. (2025) have shown strong capabilities in natural language understanding, generation, and reasoning Wei et al. (2022); Zhong et al. (2023a); Peng et al. (2023); Lu et al. (2024); Zhao et al. (2023). However, as world knowledge evolves, LLMs need methods to update outdated information efficiently. Knowledge editing Yao et al. (2023) allows modifications to specific knowledge while preserving unrelated information, making it more cost-effective than retraining from scratch.
Figure 1: Examples of (a) monolingual and (b) cross-lingual knowledge editing. In the former, the editing and verification languages are the same, while in the latter, knowledge is transferred from the source language (e.g., English) to the target language (e.g., Chinese).
Despite significant progress, most existing approaches focus on monolingual editing De Cao et al. (2021); Dai et al. (2021); Mitchell et al. (2021). As LLMs are increasingly required to handle multilingual queries Zhang et al. (2024); Wang et al. (2023b), monolingual solutions often fail. For example, when editing the response "the Conservative Party" to "the Labor Party" in English (Figure 1(a)), this update does not propagate to the Chinese version. Thus, expanding knowledge editing to cross-lingual settings is crucial to ensure that changes made in the source language are correctly applied to target languages.
Currently, several studies on multilingual knowledge editing have emerged Xu et al. (2022); Wang et al. (2023d); Wei et al. (2024); Xie et al. (2024); Liang et al. (2024). Some of these methods extend the edited language from single to multiple, while others prescribe source-language answers as the ground truth for multilingual queries. Both strategies fall short of achieving true cross-lingual knowledge democratization. For example, although IKE was regarded as the state-of-the-art method in previous studies Wang et al. (2023a); Xie et al. (2024), its performance on the Bi-ZsRE benchmark demonstrates significant limitations, achieving merely a 73.33 average score when editing in English. Unlike previous methods that attempt to forcefully correct LLM behaviors in both the source and target languages, we propose guiding LLMs to internalize knowledge from source language editing and apply it to target language queries. Our Cross-Lingual Knowledge Democracy Edit (X-KDE) with Dual-Stage Refinement, where we use parallel language datasets to transfer knowledge from the source to the target language.
The X-KDE framework involves two phases: (i) Cross-lingual edition Instruction Tuning (XE-IT), where the source language editing descriptor is paired with target language queries to create a parallel dataset, guiding the model to answer in the target language while preserving unchanged knowledge. (ii) Target-language Preference Optimization (TL-PO), where we adopt the ORPO strategy Hong et al. (2024) further constrains cross-lingual knowledge, promoting the diffusion of updates from source to target languages, and achieving true knowledge democratization. Taking the Bi-ZsRE benchmark as an example, X-KDE outperforms others, achieving average scores of 91.04 and 88.49 when editing in English and Chinese, respectively. Our contributions are three-fold:
  • To tackle the scarcity of high-quality resources in cross-lingual knowledge editing, we introduce new datasets that fill gaps in existing resources, enhancing the reliability of knowledge transfer across languages.
  • We propose X-KDE, a simple yet highly effective method for cross-lingual knowledge editing. This approach, based on a two-stage process, ensures robust knowledge generalization across languages.
  • Through extensive experiments, we establish X-KDE as a new state-of-the-art (SOTA) solution for cross-lingual knowledge editing, demonstrating significant improvements in performance while preserving original knowledge and enhancing the portability of updates.
2 Preliminary
2.1 Knowledge Editing
Knowledge editing selectively modifies in-scope knowledge while preserving out-of-scope behavior. Given a base LLM $p_{\theta}$ and an edit descriptor $\langle x_{e},y_{e}\rangle$, where $x_{e}$ is the modification description and $y_{e}$ is the corresponding answer, the edited model should adhere to four key properties:
Reliability
evaluates accuracy on edit descriptors:
$\mathbb{E}_{(x_{e},y_{e})\sim\mathcal{X}_{e}}\mathbf{1}[\mathop{\arg\max}_{y}p_{\theta}^{*}(y|x_{e})=y_{e}]$ (1)
Generality
assesses the precision of semantically rephrased examples:
$\mathbb{E}_{(x_{e}^{par},y_{e})\sim\mathcal{X}_{e}^{par}}\mathbf{1}[\mathop{\arg\max}_{y}p_{\theta}^{*}(y|x_{e}^{par})=y_{e}]$ (2)
Locality
ensures that out-of-scope inputs remain unchanged:
$\mathbb{E}_{(x_{e},y_{e})\sim\mathcal{O}_{e}}\mathbf{1}[p_{\theta}^{*}(y|x_{e})=p_{\theta}(y|x_{e})]$ (3)
Portability
measures the ability to transfer updated knowledge to related queries:
$\mathbb{E}_{(x_{e},y_{e})\sim\mathcal{I}_{e}}\mathbf{1}[\mathop{\arg\max}_{y}p_{\theta}^{*}(y|x_{e})=y_{e}]$ (4)
2.2 Cross-Lingual Knowledge Editing
Cross-lingual knowledge editing extends monolingual knowledge editing by requiring a multilingual LLM $p_{m\theta}$ to propagate knowledge from a source language to a target language. Given an edit descriptor in the source language $\langle x_{e}^{s},y_{e}^{s}\rangle$, the goal is to maximize:
$\mathbb{E}_{(x_{e}^{s},y_{e}^{s})\sim\mathcal{X}_{e}^{s}\atop x_{e}^{t}=I^{t}({x}^{s}_{e}),y_{e}^{t}=I^{t}({y}^{s}_{e})}\mathbf{1}[\mathop{\arg\max}_{y}p_{m\theta}^{*}(y|x_{e}^{t})=y_{e}^{t}]$ (5)
$\mathbb{E}_{(x_{e}^{s},y_{e}^{s})\sim\mathcal{O}_{e}^{s}\atop x_{e}^{t}=I^{t}({x}^{s}_{e}),y_{e}^{t}=I^{t}({y}^{s}_{e})}\mathbf{1}[p_{m\theta}^{*}(y|x_{t})=p_{m\theta}(y|x_{t})]$ (6)
Here, $x_{e}^{t},y_{e}^{t}$ are the edit descriptors in the target language $t$, and $I^{t}(\cdot)$ translates the source language input into the target language. Cross-lingual knowledge editing also requires cross-lingual comprehension, ensuring that updates in the source language lead to consistent responses in the target language.
3 Methodology
Figure 2: Illustration of Cross-Lingual Knowledge Democracy Edit (X-KDE) framework. In the XE-IT phase, we fine-tune the LLM on a carefully curated parallel dataset, enabling it to incorporate newly edited information from the source language when queried in the target language. In the TL-PO phase, multiple responses are generated, ranked based on similarity to the target language answer, and alignment optimization is applied to refine the output.
To achieve the democratization of knowledge, we propose the Cross-Lingual Knowledge Democracy Edit (X-KDE) framework, as shown in Figure 2. This framework enables LLMs to adapt to evolving knowledge demands and facilitates the transfer of knowledge to target languages by editing only the source language. X-KDE consists of two stages: the Cross-lingual Edition Instruction Tuning (XE-IT) phase and the Target-language Preference Optimization (TL-PO) phase.
3.1 XE-IT Stage: Learning to Edit
The goal is to enable the model to leverage knowledge edits in the source language while preserving the unchanged information. To meet the requirements for cross-lingual editing, we carefully constructed a high-quality dataset and employed XE-IT to fine-tune the model.
3.1.1 Dataset Construction
Data Sources.
Our goal is to enable the model to use edit descriptors effectively while maintaining unrelated information. We use several widely used knowledge editing datasets, including ZsRE Levy et al. (2017), HalluEditBench Huang et al. (2024a), RIPPLEEDITS Cohen et al. (2024), WikiBio Hartvigsen et al. (2024), MQUAKE Zhong et al. (2023b), and COUNTERFACT Meng et al. (2022a), to build our training data. These datasets provide edit descriptors paired with QA pairs. To mitigate data leakage, we randomly sample and translate subsets for training.
Sample Generation.
Existing datasets often feature straightforward QA pairs, which limit the model’s comprehension ability. To address this, we use Deepseek Liu et al. (2024) to generate complex in-scope and out-of-scope query-answer pairs. This method enhances training data quality and model comprehension.
Quality Control.
To ensure relevance, we use Deepseek to assess the quality of in-scope query-answer pairs. Samples are scored based on Syntactic Structure, Lexical Richness, and Edit Consistency, with low-quality samples filtered out and replaced by higher-scoring ones.
Translation Process.
We use large language models, i.e., Deepseek [^1] https://api-docs.deepseek.com/ to translate the generated data from English to Chinese.
Parallel Data Construction.
Our dataset follows a parallel structure (Figure 2(b)), with the in-scope section guiding LLMs when to use updated knowledge and the out-of-scope section minimizing the impact on unrelated knowledge. The dataset includes both monolingual and cross-lingual sections, where the source language contains edit descriptors and the target language contains queries and answers. Further details about our dataset are provided in Appendix A.3.
3.1.2 Fine-Tuning
Thanks to the flexible parallel structure, we can adaptively select the source and target languages to satisfy specific needs. We create a large-scale cross-lingual dataset and compute loss based only on the answer tokens. The model generates answers in the target language given source-language edit descriptors and target-language queries.
3.2 TL-PO Stage: Preference Optimization
3.2.1 Preference Scoring
After the XE-IT phase, the model has initially acquired the ability for cross-lingual knowledge editing. However, when faced with queries in the target language, the model may still make mistakes, such as generating responses in the source language, producing surface-level transliterations, or failing to follow target language patterns. To address this, we use a multilingual translation model to compute "alignment" scores, favoring responses aligned with the target language.
3.2.2 Alignment Optimization
We aim for the model to generate answers in the target language with higher likelihood than in the source language: $p_{\theta}^{*}(y_{e}^{t}|x_{e}^{s})>p_{\theta}^{*}(y_{e}^{s}|x_{e}^{s})$. To achieve this, we employ ORPO, a state-of-the-art preference optimization method. We collect flawed outputs ($Y_{l}$) and the ground truth in the target language ($Y_{w}$), then optimize the objective function:
$\mathcal{L}_{ORPO}=\mathbb{E}_{(x,y_{w},y_{l})}\left[\mathcal{L}_{XE-IT}+\lambda\cdot\mathcal{L}_{OR}\right]$ (7)
where $\mathcal{L}_{XE-IT}$ is the XE-IT loss and $\mathcal{L}_{OR}$ maximizes the likelihood ratio between the preferred response and the less preferred one.
$\mathcal{L}_{XE-IT}=-\frac{1}{m}\sum_{t=1}^{m}\log P_{\theta}(y_{t}|x,y_{<t})$ (8)
$\mathcal{L}_{OR}=-\log\sigma\left(\log\frac{\textbf{odds}_{\theta}(y_{w}|x)}{\textbf{odds}_{\theta}(y_{l}|x)}\right)$ (9)
$\textbf{odds}_{\theta}(y|x)=\frac{P_{\theta}(y|x)}{1-P_{\theta}(y|x)}$ (10)
@@ 4.1 Experimental Setup
Table 1: Cross-lingual editing performance of different methods, in terms of F1 score on Llama2-chat-7B backbones. Results in green indicates the best results. “Avg.” represents the overall mean of all metrics evaluated across the two languages.
@@
Table 2: Results on MzsRE dataset for editing performed in English. Here, “en-zh” means that English serves as the source language and Chinese as the target language, with similar interpretations for the other pairs. “en-avg” denotes the average performance across cross-lingual scenarios.
We chose the following methods as baselines: (1) FT-L Meng et al. (2022a) fine-tunes a specific layer of the feed-forward network to maximize the likelihood of target tokens; (2) FT-M Zhang et al. (2024) fine-tunes the same feed-forward network layer as FT-L. Additionally, it masks the original text and applies cross-entropy loss on the target answer; (3) ROME Meng et al. (2022a) employs causal mediation analysis to identify the target area for editing, and then updates the parameters of the feed-forward network layers; (4) MEMIT Meng et al. (2022c), built upon the ROME framework, enables the simultaneous update of thousands of knowledge; (5) IKE Zheng et al. (2023) utilizes the in-context-learning ability of the model and provides a few-shot demonstration to guide the model’s responses based on the updated facts. (6) LTE Jiang et al. (2024) enhances the model’s instruction-following ability through supervised fine-tuning (SFT), and employs a retrieval-based mechanism to provide updated knowledge for demonstrations.
@@
We select two public models as backbones, including LLaMA2-Chat-7B Touvron et al. (2023) and Qwen2.5-instruct-7B Yang et al. (2024a). These models are widely used in chatbot applications, the former excels in English, while the latter demonstrates strong multilingual abilities. For brevity, the results on Qwen2.5-instruct-7B are provided in D.2.
@@ 4.2 Results of Single Fact Editing
Our method is effective not only in a bilingual Chinese-English setting but can also be generalized to additional languages. We conducted more extensive experiments on the Mzsre dataset, and the results are presented in Table 2. More detailed results are in Appendix D.1. It is evident that, compared to LTE, our method exhibits 10.86% in reliability and 10.98% average gains in generality when edited in English. These results further demonstrate that our approach significantly enhances the model’s cross-lingual abilities, enabling it to effectively apply knowledge from a pivot language to others and marking a significant step toward the democratization of knowledge.
@@ 4.3 Results of Mass Fact Editing
In the previous section, we introduced the results of single-fact editing. However, real-world scenarios are often more complex, requiring simultaneous or sequential edits to multiple pieces of knowledge. Therefore, in this section, we conduct comprehensive experiments using X-KDE alongside several methods that support mass editing (FT-L, FT-M, MEMIT, and LTE) on LLaMA2-Chat-7B in both batch-edit and sequential-edit settings, and then present the corresponding results.
X-KDE can process thousands of edits simultaneously.
In line with the single-edit procedure, we evaluate both English and Chinese edits separately. For simplicity in presentation, we take the average of these two results, as shown in Figure 3. As the batch size increases, we observe a gradual decline across all performance metrics for all methods. The drop is particularly severe for MEMIT and FT-L, especially in the locality metric, which is nearly cut in half. In contrast, X-KDE achieves the best performance, maintaining the highest accuracy while exhibiting the slowest degradation rate. These results indicate that our approach remains stable in cross-lingual settings, even after thousands of edits.
X-KDE can sequentially acquire new knowledge without forgetting previous information.
In the sequential editing setting, the model integrates new knowledge on top of its previous edits, which leads to a gradual decline in performance over time. As illustrated in Figure 4, the performance of methods that modify model parameters typically degrades as the number of edits increases. For instance, MEMIT and FT-L remain stable only when the number of edits $n\leq 10$; beyond that, their performance deteriorates sharply. In contrast, knowledge storage-retrieval paradigms represented by X-KDE and LTE circumvent direct parameter modifications through external memory architectures. Moreover, X-KDE demonstrates superior cross-lingual transfer capabilities compared to LTE, achieving better performance across diverse data streams.
@@
Table 3: General tasks performance of X-KDE, and LLaMA2-Chat-7B, measured by F1 scores. Results in bold represent the best performance in each category.
@@ 4.4 Results of General Tasks
A series of studies have demonstrated that knowledge editing can influence model performance across various scenarios Yang et al. (2024b); Li et al. (2023b); Gu et al. (2024). To investigate whether our method impacts the model’s capabilities in unrelated domains, we conducted tests across a range of fields. Given that the cross-lingual knowledge editing task typically involves two languages, we use English and Chinese as representative examples. Multiple benchmarks are selected in these two languages, covering tasks such as commonsense reasoning, natural language understanding, open-domain QA, and general intelligence. For example, the benchmarks selected for English include MMLU Hendrycks et al. (2020), CommonSenseQA Talmor et al. (2018), PIQA Bisk et al. (2020), XSum Narayan et al. (2018), and Natural Questions Kwiatkowski et al. (2019). For Chinese, the chosen benchmarks are CMMLU Li et al. (2023a), CommonSenseQA_zh Contributors (2023), CEval Huang et al. (2024b), and Natural Questions_zh Contributors (2023). We conducted all experiments using the OpenCompass tool Contributors (2023). The results are presented in Table 3. Overall, our method not only preserves the model’s performance in English but also significantly enhances its capabilities in Chinese. Although certain tasks, such as CommonsenseQA, XSum, and PIQA, show a decrease in performance when tested in English, the overall results demonstrate consistent English capabilities. This highlights the robustness of our method, which achieves cross-lingual knowledge editing while preserving the model’s original performance and significantly improving its proficiency in Chinese.
@@ 5.3 Why choose ORPO as the preferred optimization method?
We evaluate several popular preference optimization methods using a held-out dataset from our training data, specifically direct policy optimization (DPO, Rafailov et al., 2023), contrastive preference optimization (CPO, Xu et al., 2024), Kahneman-Tversky Optimization (KTO, Ethayarajh et al., 2024), and odds ratio preference optimization (ORPO, Hong et al., 2024) on the Bi-ZsRE benchmark. As shown in Table 6, ORPO outperforms the other methods, achieving the best overall performance. While CPO and KTO also yield similar improvements, ORPO excels in preserving irrelevant target-language samples, demonstrating superior locality. In contrast, DPO results in a performance decline, which we attribute to the absence of negative log-likelihood (NLL) constraints, potentially weakening the model’s instruction-following capabilities. Based on these results, we adopt ORPO as the default optimization method for the second phase of our approach, as it provides the most significant improvement.
@@
6 Related Work
Knowledge Editing
The task of knowledge editing was introduced by  Sinitsin et al. (2020) to update specific knowledge while preserving unrelated information. Current methods fall into two paradigms: preserving or modifying the model’s parameters. (1) Preserving LLMs’ parameters involves auxiliary models or extra parameters. SERAC Mitchell et al. (2022) uses a counterfactual model to update knowledge without altering model parameters. TPatcher Huang et al. (2023) and CaliNET Dong et al. (2022) add trainable parameters to edit knowledge. IKE Zheng et al. (2023) and ICE Cohen et al. (2024) leverage in-context learning to correct knowledge. (2) Modifying the model’s parameters directly updates specific parameters to change knowledge. KE De Cao et al. (2021) and MEND Mitchell et al. (2021) predict weight updates for new data using a hyper-network. KN Dai et al. (2021), ROME Meng et al. (2022b), and MEMIT Meng et al. (2022c) use knowledge attribution or causal mediation analysis to target specific parameters for updating.
Cross-Lingual Knowledge Editing
Nearly all language models are multilingual, making it crucial to enhance instruction-following capabilities across different languages Zan et al. (2024) and enable synchronized cross-lingual knowledge updates Xu et al. (2022); Wang et al. (2023a). Cross-lingual knowledge editing extends monolingual editing by propagating edits across languages.  Wang et al. (2023a) introduced cross-lingual knowledge editing and created the Bi-ZsRE dataset to assess the applicability of monolingual methods in multilingual contexts. LiME Xu et al. (2022) proposes language anisotropic editing to enhance cross-lingual editing, and MPN Si et al. (2024) introduces multilingual patch neurons to update knowledge. However, these methods treat source language answers as ground truth for target language queries, falling short of achieving true cross-lingual transfer.
LLM Alignment
LLM alignment Gabriel (2020) ensures that LLMs’ behaviors align with human values. Techniques such as supervised fine-tuning (SFT) Wei et al. (2021); Wang et al. (2023e); Mishra et al. (2021); Wang et al. (2024) train models to follow task descriptions in natural language. Despite SFT, models may still generate harmful content Carlini et al. (2021); Gehman et al. (2020); Zhang et al. (2025). To address this, reinforcement learning with human feedback (RLHF) Stiennon et al. (2020); Ouyang et al. (2022) refines models further. Due to the issues of fragile training and reward hacking in RLHF Miao et al. (2024), recent simplified methods like SimPO Meng et al. (2024) and ORPO Hong et al. (2024) effectively enable preference optimization.
7 Conclusion
In this paper, we present the Cross-Lingual Knowledge Democracy Edit (X-KDE) framework, which facilitates knowledge editing across languages in large language models (LLMs). By integrating Cross-lingual Edition Instruction Tuning (XE-IT) and Target-language Preference Optimization (TL-PO), X-KDE efficiently transfers knowledge from a source language to a target language while maintaining strong performance in monolingual settings. Additionally, we introduce high-quality datasets specifically designed for cross-lingual knowledge editing, filling gaps in existing resources. Our experimental results demonstrate that X-KDE outperforms current methods, offering a scalable solution for cross-lingual knowledge editing. Future research will explore applying X-KDE to other domains and optimizing its efficiency.
Limitations
While our work presents promising results, there are a few limitations to consider. First, due to computational constraints, we validate X-KDE on models with up to 7B parameters. Evaluating larger models, such as those exceeding 70B parameters, could provide more robust insights. Second, while our method has been effective in multilingual settings, its application to additional domains, such as finance or law, remains unexplored. Future research will focus on scaling X-KDE and extending its applicability to other fields.

### v2

@@ Abstract
Knowledge editing allows for efficient adaptation of large language models (LLMs) to new information or corrections without requiring full retraining. However, prior methods typically focus on either single-language editing or basic multilingual editing, failing to achieve true cross-linguistic knowledge synchronization. To address this, we present a simple and practical state-of-the-art (SOTA) recipe Cross-Lingual Knowledge Democracy Edit (X-KDE), designed to propagate knowledge from a dominant language to other languages effectively. Our X-KDE comprises two stages: (i) Cross-lingual Edition Instruction Tuning (XE-IT), which fine-tunes the model on a curated parallel dataset to modify in-scope knowledge while preserving unrelated information, and (ii) Target-language Preference Optimization (TL-PO), which applies advanced optimization techniques to ensure consistency across languages, fostering the transfer of updates. Additionally, we contribute a high-quality, cross-lingual dataset, specifically designed to enhance knowledge transfer across languages. Extensive experiments on the Bi-ZsRE and MzsRE benchmarks show that X-KDE significantly enhances cross-lingual performance, achieving an average improvement of +8.19%, while maintaining high accuracy in monolingual settings. The code and dataset will be available upon acceptance. Our code is available at: ahttps://github.com/YukinoshitaKaren/X_KDE.
@@
Anonymous ACL submission
@@ 1.1 Experimental Setup
Table 1: Cross-lingual editing performance of different methods on Llama2-chat-7B backbones. Results in green indicates the best results. “Avg.” represents the overall mean of all metrics evaluated across the two languages.
@@
We chose the following methods as baselines: (1) FT-L meng2022locating fine-tunes a specific layer of the feed-forward network to maximize the likelihood of target tokens; (2) FT-M zhang2024comprehensive fine-tunes the same feed-forward network layer as FT-L. Additionally, it masks the original text and applies cross-entropy loss on the target answer; (3) ROME meng2022locating employs causal mediation analysis to identify the target area for editing, and then updates the parameters of the feed-forward network layers; (4) MEMIT meng2022mass, built upon the ROME framework, enables the simultaneous update of thousands of knowledge; (5) IKE zheng2023can utilizes the in-context-learning ability of the model and provides a few-shot demonstration to guide the model’s responses based on the updated facts. (6) LTE jiang2024learning enhances the model’s instruction-following ability through supervised fine-tuning (SFT), and employs a retrieval-based mechanism to provide updated knowledge for demonstrations.
Dataset.
We evaluate X-KDE and Baselines using two widely used benchmarks: Bi-ZsRE dataset wang2023cross and MzSRE dataset wang2023retrieval. The former contains 1,037 test samples in English and an equal number in Chinese, while the latter covers twelve languages: English, Czech, German, Dutch, Spanish, French, Portuguese, Russian, Thai, Turkish, Vietnamese and Chinese, with 743 samples in each.
@@
We select two public models as backbones, including LLaMA2-Chat-7B touvron2023llama and Qwen2.5-instruct-7B yang2024qwen2. These models are widely used in chatbot applications, the former excels in English, while the latter demonstrates strong multilingual abilities. For brevity, the results on Qwen2.5-instruct-7B are provided in D.2.
@@ 1.2 Results of Single Fact Editing
Our method is effective not only in a bilingual Chinese-English setting but can also be generalized to additional languages. We conducted more extensive experiments on the Mzsre dataset, and the results are presented in Table 2. More detailed results in Appendix D.1. It is evident that, compared to LTE, our method exhibits 10.86% in reliability and 10.98% average gains in generality when edit in English. These results further demonstrate that, our approach significantly enhances the model’s cross-lingual abilities, enabling it to effectively apply knowledge from a pivot language to others and marking a significant step toward the democratization of knowledge.
@@ 1.3 Results of Mass Fact Editing
In the previous section, we introduced the results of single fact editing. However, real-world scenarios are often more complex, requiring simultaneous or sequential edits to multiple pieces of knowledge. Therefore, in this section, we conduct comprehensive experiments using X-KDE alongside several methods that support mass editing (FT-L, FT-M, MEMIT, and LTE) on LLaMA2-Chat-7B in both batch-edit and sequential-edit settings, and then present the corresponding results.
@@
Table 3: General tasks performance of X-KDE, and LLaMA2-Chat-7B. Results in bold represent the best performance in each category.
X-KDE can process thousands of edits simultaneously.
In line with the single-edit procedure, we evaluate both English and Chinese edits separately. For simplicity in presentation, we take the average of these two results, as shown in Figure 1. As the batch size increases, we observe a gradual decline across all performance metrics for all method. The drop is particularly severe for MEMIT and FT-L, especially in the locality metric, which is nearly cut in half. In contrast, X-KDE achieves the best performance, maintaining the highest accuracy while exhibiting the slowest degradation rate. These results indicate that our approach remains stable in cross-lingual settings, even after thousands of edits.
X-KDE can sequential acquire new knowledge without forgetting previous information.
In the sequential-editing setting, the model integrates new knowledge on top of its previous edits, which leads to a gradual decline in performance over time. As illustrated in Figure 2, the performance of methods that modify model parameters typically degrades as the number of edits increases. For instance, MEMIT and FT-L remain stable only when the number of edits $n\leq 10$; beyond that, their performance deteriorates sharply. In contrast, knowledge storage-retrieval paradigms represented by X-KDE and LTE circumvent direct parameter modifications through external memory architectures. Moreover, X-KDE demonstrates superior cross-lingual transfer capabilities compared to LTE, achieving better performance across diverse data streams.
@@ 1.4 Results of General Tasks
A series of studies have demonstrated that knowledge editing can influence model performance across various scenarios yang2024butterfly; li2023unveiling; gu2024model. To investigate whether our method impact the model’s capabilities in unrelated domains, we conducted tests across a range of fields. Given that the cross-lingual knowledge editing task typically involves two languages, we use English and Chinese as representative examples. Multiple benchmarks are selected in these two languages, covering tasks such as commonsense reasoning, natural language understanding, open-domain QA, and general intelligence. For example, the benchmarks selected for English include MMLU hendrycks2020measuring, CommonSenseQA talmor2018commonsenseqa, PIQA bisk2020piqa, XSum narayan2018don, and Natural Questions kwiatkowski2019natural. For Chinese, the chosen benchmarks are CMMLU li2023cmmlu, CommonSenseQA_zh 2023opencompass, CEval huang2024c, and Natural Questions_zh 2023opencompass. We conducted all experiments using the OpenCompass tool 2023opencompass. The results are presented in Table 3. Overall, our method not only preserves the model’s performance in English but also significantly enhances its capabilities in Chinese. Although certain tasks, such as CommonsenseQA, XSum, and PIQA, show a decrease in performance when tested in English, the overall results demonstrate consistent English capabilities. This highlights the robustness of our method, which achieves cross-lingual knowledge editing while preserving the model’s original performance and significantly improving its proficiency in Chinese.
@@ 2.3 Why choose ORPO as the preferred optimization method?
We evaluate several popular preference optimization methods using a held-out dataset from our training data, specifically direct policy optimization (DPO)rafailov2023direct, contrastive preference optimization (CPO)xu2024contrastive, Kahneman-Tversky Optimization (KTO)ethayarajh2024kto, and odds ratio preference optimization (ORPO)hong2403orpo on the Bi-ZsRE benchmark. As shown in Table 6, ORPO outperforms the other methods, achieving the best overall performance. While CPO and KTO also yield similar improvements, ORPO excels in preserving irrelevant target-language samples, demonstrating superior locality. In contrast, DPO results in a performance decline, which we attribute to the absence of negative log-likelihood (NLL) constraints, potentially weakening the model’s instruction-following capabilities. Based on these results, we adopt ORPO as the default optimization method for the second phase of our approach, as it provides the most significant improvement.